OpenStudy (anonymous):

Solve p(x)=x^4 +6x^2 +9

OpenStudy (anonymous):

There's nothing to solve there.

OpenStudy (anonymous):

You probably mean solve p(x) = 0

OpenStudy (anonymous):

isnt this where you do the (\[(x ^{2}+3)(x ^{2}+3)\]

OpenStudy (anonymous):

In these you always define a variable to be x^2 and rewrite the function to be quadratic.

OpenStudy (anonymous):

p(x) = x^4 +6x^2 +9 = 0 t^2 + 6t + 9 = 0, t=x^2 solve like you would a quadratic equation.

OpenStudy (anonymous):

\[x^4+6 x^2+9=\left(x^2+3\right)^2 \] @arhendrickson Were you asking that your problem expression be factored in the first place?